Any two non-zero subgroups of Z are isomorphic. 


1
Expert's answer
2021-07-19T05:51:19-0400

Solution:

Yes, it is correct. Let "H \leq \mathbb{Z}" be a subgroup. If "H=\{0\}" then "H=0 \mathbb{Z}" and we are done. Otherwise H must contain non-zero elements, some of them positive. Let "n=\min \{k>0: k \in H\}" . Then it can be shown that "H=n \mathbb{Z}" . We have "n \in H" and hence obviously "n \mathbb{Z} \leq H" , because H is closed to addition and additive inverses. For the other direction let "k \in H" . We can divide k by n with remainder. We get an expression of the form k=q n+r where "q, r \in \mathbb{Z}, 0 \leq r<n" . But then note that "r=k+(-q) n \in H" . Since n is by definition the smallest positive number in H and "0 \leq r<n" we have to conclude that r=0, and hence "k \in n \mathbb{Z}" . This shows that "H \leq n \mathbb{Z}" .

If "n \neq 0\ then\ n \mathbb{Z}" is indeed isomorphic to "\mathbb{Z}" . We can define an isomorphism "\varphi: \mathbb{Z} \rightarrow n \mathbb{Z}\ by\ k \rightarrow n k"